-- Main.jill - 13 Jan 2007

-- Main.gredel - 15 Jan 2007

afternoon, no wind no clouds no ice - sorry, lots of ice on the roof. Snow melting heaters not very efficient, build up of insulating gap between ice and roof metal which prevents efficient melting.

no effect of electric heaters over 1/4 of roof, air temp from right front heaters 15 C compared to 35 C for rest; EH1004 and EH1006 not working, J. Little being contacted for assistance

in the meanwhile, manual removal & breaking of ice

23:00 init LBC --> ERROR, Blue Channel failed, ROTATOR bad communication with serial device --> reboot cmu, OK

00:23 25x bias (U-band) --> 880 ADU

00:40 opening OPEN AT LAST

00:53 Skyflat U-band 1sec 5x dithered --> saturated

01:01 repeat sky U 1 sec dither 5 --> 21000 ADU 01:06 sky U-band 1 sec Rotator 180 dither 5 --> 13000 ADU

Temperatures: mirror -5 C, ambient -12 C

01:10 sky B 2 sec 5 dither --> 17000 ADU 01:15 sky B 5 sec 5 dither --> 20000 ADU 01:18 sky B 10 sec RA 180 5 dither --> 16000 ADU few stars visible, way out of focus 01:24 sky V 30 sec 5 dither --> 9500 ADU 01:31 sky V 60 sec RA 180 13000 01:42 Image Analysis lbcBfastextra
Computed Aberrations (nm) Defoc AstX AstY ComaX ComaY Sph (Z4) (Z5) (Z6) (Z7) (Z8) (Z11)
429 -193 377 999 2526 -263 383 1062 -692 -586 -766 389 -6943 0 0 0 0 0 -540 265 -32 667 173 -59 186 -212 942 -686 -337 40 -186 -240 -374 846 516 253

02:27 take image 1 x 10 sec B --> 1.6 arcsec seeing 02:29 L95_BV standard field, 1 and 2 sec in B and V IQ again, image pupil does not look very good clear all, add 1000 sperical

-1087 344 131 -709 -155 -342

02:49 Standard field again L_95 U 1, 4, 10 sec --> 2'' 03:03 PG0231 U, B, V 03:07 L92 U, B, V seeing about 1.5'', temperature difference mirror - dome > 4 C IA again preset to Ryan 0deg_jan_epoch1, ROTATOR FAILURE power off/on , reboot etc. failure, error message 'rotator moves by 0.4 degrees instead of 56 degrees'

phone interactions with John Hill --> NO ELEC or other TEC support on the mountain

2nd problem occurs with pressure indicators in mirror cell, interactions Jonathan - Ed (it is good and efficient that Jonathan is up on the mountain)

OB execution errors -- felft main mirror failed to focus

enter technical rotator control program, rotator back to work after serveral manual movements back and forth

06:43 preset to Kochanek M81 not clear if IA works well, not clear if poor outside seeing or spherical -- when do we get a DIMM seeing B 2 - 2.5 arcsec 07:08 Kochanek M81 BV but poor seeing, possibly not useful, will try again IA at higher elevation

07:30 IA at 67 near Ryan 0deg asteroid, very slow, LBC control interface seems to hang IA does not seem to be OK, reset all parameters, IA again 07:54 Ryan 0deg_jan_epoch 1 --> looks like the first useful science observation preset to M81 --> ROTATOR FAILURE

09:06 Ryan 0deg_jan_epoch2 preset M81 --. ROTATOR AGAIN --> sorry Chris, will need to check this before going back to M81 09:49 Ryan 0deg_jan_epoch3

* Ryan 0deg three epochs Jan complete *

preset NGC 5907 Fan ROTATOR FAILURE

no further objects in poor seeing, frequent rotator failures, not sure about IA

good: mirror cell seems to work fine at -12 bad: mirror takes very long to cool, mirror was 5 deg too warm various technical problems, no tec support on site (except of JKern, and that worked well)

CLOSE
Topic revision: r4 - 16 Jan 2007, gredel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback